Air is compressed in a cylinder of a diesel engine. The initial volume of the air is 800cm3 and a temperature of 27oC and a pressure of 100kPa. The air compressed to one-fifteenth of its initial volume so that its temperature reaches 600oC. Determine the final pressure of the air.


1
Expert's answer
2021-06-02T13:12:02-0400

Gives

"T_1=27\u00b0C=300K"

"P_1=10^5pa"

"T_2=600\u00b0C=873K"

"V_1=8\times10^{-4}m^3"

"V_2=\frac{V_1}{15}"

"P_2=?"

Gas equation

PV= nRT

n=1mol

"P_2=\frac{P_1V_1}{V_2}\times\frac{T_2}{T_1}"

Put values


"P_2=\frac{10^5\times V_1}{\frac{V_1}{15}}\times\frac{873}{300}=4365Kpa"

"P_2=4365 Kpa"